OpenStudy (anonymous):

How do I find the derivative of 2sinxcosx

OpenStudy (anonymous):

you have a choice you can use the product rule of you can recall some trig identity

OpenStudy (anonymous):

product rule sounds awesome

OpenStudy (anonymous):

can you tell me what i'm doing wrong first?

OpenStudy (anonymous):

second way is probably easiest, \[2\sin(x)\cos(x)=\sin(2x)\]

OpenStudy (anonymous):

i don't know that identity....

OpenStudy (anonymous):

if I do 2sinxcosx then dy/dx = 2((sinx)(-sinx) + (cosx)(cosx))

OpenStudy (anonymous):

ok go ahead and i will check, but if the answer in the back of the book is something different, it is probably not that you did it wrong, but rather than they used some identity you don't know (aka remember)

OpenStudy (anonymous):

it's not in the back of the book, i'm just pretty positive i don't understand this

OpenStudy (anonymous):

no you have it correct \[2(-\sin^2(x)+\cos^2(x))\]

OpenStudy (anonymous):

but is is also \[2\cos(2x)\]

OpenStudy (anonymous):

i mean "yes, your answer is correct"

OpenStudy (anonymous):

alright, so then an identity of cos^2x is 1-sin^2x right?

OpenStudy (anonymous):

why is it 2cos2x haha

OpenStudy (anonymous):

the easiest reason i can think of is that \[\sin(2x)=2\sin(x)\cos(x)\]

OpenStudy (anonymous):

i don't understand that though..

OpenStudy (anonymous):

you mean why it is true?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

depends on where you are willing to start, and what you know

OpenStudy (anonymous):

if you grant that \[\sin(\alpha +\beta)=\sin(\alpha)\cos(\beta)+\sin(\beta)\cos(\alpha)\] the replace both \(\alpha\) and \(\beta\) by \(x\) and you get it in one step

OpenStudy (anonymous):

if not, you have to use a bunch of tedious geometry to get the "addition angle" formula, and then replace \(\alpha\) and \(\beta\) by \(x\)

OpenStudy (anonymous):

i've never heard of the addition angle. i know the product rule though

OpenStudy (anonymous):

wait that's not the product rule

OpenStudy (anonymous):

i know neither, i give up

OpenStudy (anonymous):

well you already showed that you know the product rule

OpenStudy (anonymous):

trig identities have nothing to do with calculus, they are just identities there is a nice list here http://math2.org/math/trig/identities.htm i wouldn't bother memorizing them just refer to a list when you need them

OpenStudy (anonymous):

but you do see on that list if you look down a little that \[\cos(2x)=\cos^2(x)-\sin^2(x)\] so your derivative, which was \[2(\cos^2(x)-\sin^2(x))\] can be rewritten as \[2\cos(2x)\] if you like

OpenStudy (anonymous):

i'll keep that in mind, thank you!